OpenStudy (anonymous):

The average (arithmetic mean) of 5,10,15 and x is 20..What is the x ? How do I solve this question?

OpenStudy (littlebird):

\[\frac{ 5+10+15 +x}{ 4 }\]=20

OpenStudy (littlebird):

\[\frac{ 30+x }{ 4 }=20\]

OpenStudy (littlebird):

30+x=80

OpenStudy (littlebird):

X=50

OpenStudy (littlebird):

Another way you could have done it....

OpenStudy (littlebird):

20x4=80

OpenStudy (littlebird):

80-15-10-5=50

OpenStudy (anonymous):

Got it! Thanks.